### v1

@@ 5.2.1 Proof of Theorem 3
Proof of Theorem 3  Suppose there are infinite users and $\mathcal{D}$ be the uniform distribution. For each user $x$, there is an individual set of $12$ nodes, denoted as $N(x)=\{v_{x,1},v_{x,2},\ldots,v_{x,12}\}$. The graph $g_{1}$’s neighborhood is the first $8$ nodes $N_{g_{1}}(x)=\{v_{x,1},v_{x,2},\ldots,v_{x,8}\}$ and the graph $g_{2}$’s neighborhood is the last $8$ nodes $N_{g_{2}}(x)=\{v_{x,5},v_{x,6},\ldots,v_{x,12}\}$. There are two worlds in which the target graph $g^{\text{target}}$ is generated differently:
  • World I: W.p. $\frac{1}{2}$, $N_{g^{\text{target}}}(x)=N_{g_{1}}(x)$; w.p. $\frac{1}{2}$, $N_{g^{\text{target}}}(x)=\{u_{1},u_{2}\}$ where $u_{1}$ is sampled uniformly from $\{v_{x,5},v_{x,6},\ldots,v_{x,8}\}$ and $u_{2}$ is sampled uniformly from $\{v_{x,9},v_{x,10},\ldots,v_{x,12}\}$.
  • World II: W.p. $\frac{1}{2}$, $N_{g^{\text{target}}}(x)=N_{g_{2}}(x)$; w.p. $\frac{1}{2}$, $N_{g^{\text{target}}}(x)=\{u_{1},u_{2}\}$ where $u_{1}$ is sampled uniformly from $\{v_{x,5},v_{x,6},\ldots,v_{x,8}\}$ and $u_{2}$ is sampled uniformly from $\{v_{x,1},v_{x,2},\ldots,v_{x,4}\}$.
@@
In both worlds, the distribution of $v$ is identical, i.e., w.p. $\frac{1}{2}$, $v$ is sampled from $\mathrm{Unif}(N_{g_{1}}(x))$ and w.p. $\frac{1}{2}$, $v$ is sampled from $\mathrm{Unif}(N_{g_{2}}(x))$. Hence, if no $x$ has been sampled twice, no algorithm can distinguish the two worlds. Since there are infinite users, we will not sample the same user twice almost surely. For any output graph $g^{\text{output}}$, we let
$n_{1}=|N_{g^{\text{output}}}(x)\cap\{v_{x,1},v_{x,2},\ldots,v_{x,4}\}|\in\{0,1,\ldots,4\}\,,$
$n_{2}=|N_{g^{\text{output}}}(x)\cap\{v_{x,5},v_{x,6},\ldots,v_{x,8}\}|\in\{0,1,\ldots,4\}\,,$
$n_{3}=|N_{g^{\text{output}}}(x)\cap\{v_{x,9},v_{x,10},\ldots,v_{x,12}\}|\in\{0,1,\ldots,4\}\,.$
@@ 5.2.2 Proof of Theorem 1
Proof of Theorem 1  Let $\Delta=\sqrt{\frac{\log(|\mathcal{H}|/\delta)}{m}}$, $\overline{r}=r+2\Delta$ and $\overline{p}=p+2\Delta$. Let $g^{\dagger}$ denote the graph with precision and recall losses $(p,r)$. We know that with probability at least $1-\delta$, all graphs in $\widehat{\mathcal{H}}$ have empirical recall loss no greater than $\overline{r}$ and $g^{\dagger}$ also has ^$\widehat{\ell}^{\textrm{precision}}(g^{\dagger})\leq\overline{p}$. Then the proof is divided into two parts:
@@
    graph $g^{\dagger}$ satisfies that
@@
    any graph $g^{\prime}$ satisfying $\max_{g\in\widehat{\mathcal{H}}}\sum_{i\in[m]}\log\frac{n_{g^{\prime}}(x_{i})\wedge 4n_{g}(x_{i})}{n_{g}(x_{i})\wedge 4n_{g^{\prime}}(x_{i})}\leq 6m\overline{r}+4m\overline{p}+2$ has ^$\widehat{\ell}^{\textrm{precision}}(g^{\prime})\leq 28r+15p+o(1)$. Hence, $g^{\text{output}}$ can achieve good precision.
@@
For any $x_{1},\ldots,x_{m}$ and graph $g$ with ^$\widehat{\ell}^{\textrm{recall}}(g)\leq\overline{r}$, we have
@@
Proof  Let $E=\{i|\frac{n_{g}(x_{i})}{n_{g^{\text{target}}}(x_{i})}\geq\frac{1}{2},\frac{1}{2}\leq\frac{n_{g^{\dagger}}(x_{i})}{n_{g^{\text{target}}}(x_{i})}\leq 2\}$. According to Lemma 6, we know |$|\neg E|\leq 4m\overline{r}+2m\overline{p}$.
@@
For any $x_{1},\ldots,x_{m}$, any graph $g$, and any positive constant $c\in(0,1]$, the empirical loss for precision ^$\widehat{\ell}^{\textrm{precision}}(g)$ satisfies
^$\widehat{\ell}^{\textrm{precision}}(g)\leq\frac{1+c}{m}\sum_{i\in[m]:\frac{n_{g}(x_{i})}{n_{g^{\text{target}}}(x_{i})}\geq 1}\log\frac{n_{g}(x_{i})\wedge 2n_{g^{\text{target}}}(x_{i})}{n_{g^{\text{target}}}(x_{i})}+\frac{1+c}{c}\widehat{\ell}^{\textrm{recall}}(g)\,.$
@@
^$\widehat{\ell}^{\textrm{precision}}(g)$
$=$ $\frac{1}{m}\sum_{i=1}^{m}\frac{|N_{g}(x_{i})\setminus N_{g^{\text{target}}}(x_{i})|}{n_{g}(x_{i})}$
$\leq$ $\frac{1}{m}\sum_{i\in A_{g}}\frac{|N_{g}(x_{i})\setminus N_{g^{\text{target}}}(x_{i})|}{n_{g}(x_{i})}+\frac{1}{m}\sum_{i=1}^{m}\mathds{1}(i\notin A_{g})$
$\leq$ $\frac{1}{m}\sum_{i\in A_{g}}\min\left(\frac{(1+c)|N_{g}(x_{i})\setminus N_{g^{\text{target}}}(x_{i})|}{n_{g^{\text{target}}}(x_{i})},1\right)+\frac{1+c}{c\cdot m}\sum_{i\notin A_{g}}\frac{|N_{g^{\text{target}}}(x_{i})\setminus N_{g}(x_{i})|}{n_{g^{\text{target}}}(x_{i})}$
$\leq$ $\frac{1+c}{m}\sum_{i\in A_{g}}\min\left(\frac{n_{g}(x_{i})+|N_{g^{\text{target}}}(x_{i})\setminus N_{g}(x_{i})|-n_{g^{\text{target}}}(x_{i})}{n_{g^{\text{target}}}(x_{i})},1\right)+\frac{1+c}{c\cdot m}\sum_{i\notin A_{g}}\frac{|N_{g^{\text{target}}}(x_{i})\setminus N_{g}(x_{i})|}{n_{g^{\text{target}}}(x_{i})}$
$\leq$ $\frac{1+c}{m}\sum_{i\in A_{g}}\min\left(\frac{n_{g}(x_{i})}{n_{g^{\text{target}}}(x_{i})}-1,1\right)+\frac{1+c}{m}\sum_{i\in A_{g}}\frac{|N_{g^{\text{target}}}(x_{i})\setminus N_{g}(x_{i})|}{n_{g^{\text{target}}}(x_{i})}+\frac{1+c}{c\cdot m}\sum_{i\notin A_{g}}\frac{|N_{g^{\text{target}}}(x_{i})\setminus N_{g}(x_{i})|}{n_{g^{\text{target}}}(x_{i})}$
@@
For any $g$ satisfying ^$\widehat{\ell}^{\textrm{recall}}(g)\leq\overline{r}$ and $\frac{1}{m}\sum_{i\in[m]}\log\frac{n_{g}(x_{i})\wedge 4n_{g^{\dagger}}(x_{i})}{n_{g^{\dagger}}(x_{i})\wedge 4n_{g}(x_{i})}\leq 6\overline{r}+4\overline{p}+\frac{2}{m}$, we have $\frac{1}{m}\sum_{i\in[m]:\frac{n_{g}(x_{i})}{n_{g^{\text{target}}}(x_{i})}\geq 1}\log\frac{n_{g}(x_{i})\wedge 2n_{g^{\text{target}}}(x_{i})}{n_{g^{\text{target}}}(x_{i})}\leq 18\overline{r}+12\overline{p}+\frac{4}{m}$.
Proof of Lemma 3  Let $B=\{i|\frac{n_{g}(x_{i})}{n_{g^{\text{target}}}(x_{i})}\geq\frac{1}{2}\}$ and $C=\{i|\frac{1}{2}\leq\frac{n_{g^{\dagger}}(x_{i})}{n_{g^{\text{target}}}(x_{i})}\leq 2\}$. For any $i\in B$, the value of $\log\frac{n_{g}(x_{i})\wedge 2n_{g^{\text{target}}}(x_{i})}{n_{g^{\text{target}}}(x_{i})}$ is in $[-1,1]$. Then we have
@@
$=$ $\sum_{i\in C\cap B}\log\frac{n_{g}(x_{i})\wedge 4n_{g^{\dagger}}(x_{i})}{n_{g^{\dagger}}(x_{i})\wedge 4n_{g}(x_{i})}+\sum_{i\in C\cap B}\log\frac{n_{g^{\dagger}}(x_{i})\wedge 4n_{g}(x_{i})}{n_{g^{\text{target}}}(x_{i})}+|\neg C|-\sum_{i\in B}\log\frac{n_{g}(x_{i})\wedge n_{g^{\text{target}}}(x_{i})}{n_{g^{\text{target}}}(x_{i})}$
$\leq$ $\sum_{i=1}^{m}\log\frac{n_{g}(x_{i})\wedge 4n_{g^{\dagger}}(x_{i})}{n_{g^{\dagger}}(x_{i})\wedge 4n_{g}(x_{i})}+2|\neg C|+2|\neg B|+\sum_{i\in C\cap B}\log\frac{n_{g^{\dagger}}(x_{i})}{n_{g^{\text{target}}}(x_{i})}+|\neg C|-\sum_{i\in B}\log\frac{n_{g}(x_{i})\wedge n_{g^{\text{target}}}(x_{i})}{n_{g^{\text{target}}}(x_{i})}$
@@
For any graph $g$ with ^$\widehat{\ell}^{\textrm{recall}}(g)\leq\overline{r}$ and any subset $S\subset[m]$, we have
$\sum_{i\in S\cap B}\log\frac{n_{g}(x_{i})\wedge n_{g^{\text{target}}}}{n_{g^{\text{target}}}(x_{i})}\geq-2m\overline{r}-1\,.$
@@
$\min\sum_{i\in S\cap B}\log\frac{n_{g}(x_{i})\wedge n_{g^{\text{target}}}}{n_{g^{\text{target}}}(x_{i})}$
@@
$\min\sum_{i\in S\cap B}\log\frac{n_{g}(x_{i})\wedge n_{g^{\text{target}}}}{n_{g^{\text{target}}}(x_{i})}\geq-2m\overline{r}-1\,.$
@@
$\sum_{i\in S}\log\frac{n_{g}(x_{i})\wedge n_{g^{\text{target}}}}{n_{g^{\text{target}}}(x_{i})}\geq\min\sum_{i\in S\cap B}\log\frac{n_{g}(x_{i})\wedge n_{g^{\text{target}}}}{n_{g^{\text{target}}}(x_{i})}-|\neg B|\geq-4m\overline{r}-1\,.$
@@
For any graph $g$ with ^$\widehat{\ell}^{\textrm{precision}}(g)\leq\overline{p}$ and any subset $S\subset[m]$, we have
@@
For any $g$ withe ^$\widehat{\ell}^{\textrm{recall}}(g)\leq\overline{r}$, we have
@@
For any $g$ withe ^$\widehat{\ell}^{\textrm{precision}}(g)\leq\overline{p}$, we have
@@
Proof  When $\frac{n_{g}(x_{i})}{n_{g^{\text{target}}}(x_{i})}<\frac{1}{2}$, we have

### v2

@@ 5.2.1 Proof of Theorem 3
Proof of Theorem 3 Consider the size of the input space $\mathcal{X}$ being infinite and $\mathcal{D}$ being the uniform distribution over it. For each input $x$, there is an individual set of $12$ items, denoted as $N(x)=\{v_{x,1},v_{x,2},\ldots,v_{x,12}\}$. The hypothesis $g_{1}$ outputs the first $8$ items $g_{1}(x)=\{v_{x,1},v_{x,2},\ldots,v_{x,8}\}$ and the hypothesis $g_{2}$ outputs the last $8$ items $g_{2}(x)=\{v_{x,5},v_{x,6},\ldots,v_{x,12}\}$. There are two worlds in which the target hypothesis $g^{\text{target}}$ is generated differently:
  • World I: W.p. $\frac{1}{2}$, $g^{\text{target}}(x)=g_{1}(x)$; w.p. $\frac{1}{2}$, $g^{\text{target}}(x)=\{u_{1},u_{2}\}$ where $u_{1}$ is sampled uniformly from $\{v_{x,5},v_{x,6},\ldots,v_{x,8}\}$ and $u_{2}$ is sampled uniformly from $\{v_{x,9},v_{x,10},\ldots,v_{x,12}\}$.
  • World II: W.p. $\frac{1}{2}$, $g^{\text{target}}(x)=g_{2}(x)$; w.p. $\frac{1}{2}$, $g^{\text{target}}(x)=\{u_{1},u_{2}\}$ where $u_{1}$ is sampled uniformly from $\{v_{x,5},v_{x,6},\ldots,v_{x,8}\}$ and $u_{2}$ is sampled uniformly from $\{v_{x,1},v_{x,2},\ldots,v_{x,4}\}$.
@@
In both worlds, the distribution of $v$ is identical, i.e., w.p. $\frac{1}{2}$, $v$ is sampled from $\mathrm{Unif}(g_{1}(x))$ and w.p. $\frac{1}{2}$, $v$ is sampled from $\mathrm{Unif}(g_{2}(x))$. Hence, if no $x$ has been sampled twice, no algorithm can distinguish the two worlds. Since the input space is huge, we will not sample the same input twice almost surely. For any output hypothesis $g^{\text{output}}$, we let
$n_{1}=|g^{\text{output}}(x)\cap\{v_{x,1},v_{x,2},\ldots,v_{x,4}\}|\in\{0,1,\ldots,4\}\,$,
$n_{2}=|g^{\text{output}}(x)\cap\{v_{x,5},v_{x,6},\ldots,v_{x,8}\}|\in\{0,1,\ldots,4\}\,$,
$n_{3}=|g^{\text{output}}(x)\cap\{v_{x,9},v_{x,10},\ldots,v_{x,12}\}|\in\{0,1,\ldots,4\}\,.$
@@ 5.2.2 Proof of Theorem 1
Proof of Theorem 1 Let $\Delta=\sqrt{\frac{\log(|\mathcal{H}|/\delta)}{m}}$, $\overline{r}=r+2\Delta$ and $\overline{p}=p+2\Delta$. Let $g^{\dagger}$ denote the hypothesis with precision and recall losses $(p,r)$. We know that with probability at least $1-\delta$, all hypotheses in $\widehat{\mathcal{H}}$ have empirical recall loss no greater than $\overline{r}$ and $g^{\dagger}$ also has $\widehat{\ell}^{\textrm{precision}}(g^{\dagger})\leq\overline{p}$. Then the proof is divided into two parts:
@@
    hypothesis $g^{\dagger}$ satisfies that
@@
    any hypothesis $g^{\prime}$ satisfying $\max_{g\in\widehat{\mathcal{H}}}\sum_{i\in[m]}\log\frac{n_{g^{\prime}}(x_{i})\wedge 4n_{g}(x_{i})}{n_{g}(x_{i})\wedge 4n_{g^{\prime}}(x_{i})}\leq 6m\overline{r}+4m\overline{p}+2$ has $\widehat{\ell}^{\textrm{precision}}(g^{\prime})\leq 28r+15p+o(1)$. Hence, $g^{\text{output}}$ can achieve good precision.
@@
For any $x_{1},\ldots,x_{m}$ and hypothesis $g$ with $\widehat{\ell}^{\textrm{recall}}(g)\leq\overline{r}$, we have
@@
Proof Let $E=\{i|\frac{n_{g}(x_{i})}{n_{g^{\text{target}}}(x_{i})}\geq\frac{1}{2},\frac{1}{2}\leq\frac{n_{g^{\dagger}}(x_{i})}{n_{g^{\text{target}}}(x_{i})}\leq 2\}$. According to Lemma 6, we know $|\neg E|\leq 4m\overline{r}+2m\overline{p}$.
@@
For any $x_{1},\ldots,x_{m}$, any hypothesis $g$, and any positive constant $c\in(0,1]$, the empirical loss for precision $\widehat{\ell}^{\textrm{precision}}(g)$ satisfies
$\widehat{\ell}^{\textrm{precision}}(g)\leq\frac{1+c}{m}\sum_{i\in[m]:\frac{n_{g}(x_{i})}{n_{g^{\text{target}}}(x_{i})}\geq 1}\log\frac{n_{g}(x_{i})\wedge 2n_{g^{\text{target}}}(x_{i})}{n_{g^{\text{target}}}(x_{i})}+\frac{1+c}{c}\widehat{\ell}^{\textrm{recall}}(g)\,.$
@@
$\widehat{\ell}^{\textrm{precision}}(g)$
$=$ $\frac{1}{m}\sum_{i=1}^{m}\frac{|g(x_{i})\setminus{g^{\text{target}}}(x_{i})|}{n_{g}(x_{i})}$
$\leq$ $\frac{1}{m}\sum_{i\in A_{g}}\frac{|g(x_{i})\setminus g^{\text{target}}(x_{i})|}{n_{g}(x_{i})}+\frac{1}{m}\sum_{i=1}^{m}\mathds{1}(i\notin A_{g})$
$\leq$ $\frac{1}{m}\sum_{i\in A_{g}}\min\left(\frac{(1+c)|g(x_{i})\setminus g^{\text{target}}(x_{i})|}{n_{g^{\text{target}}}(x_{i})},1\right)+\frac{1+c}{c\cdot m}\sum_{i\notin A_{g}}\frac{|g^{\text{target}}(x_{i})\setminus g(x_{i})|}{n_{g^{\text{target}}}(x_{i})}$
$\leq$ $\frac{1+c}{m}\sum_{i\in A_{g}}\min\left(\frac{n_{g}(x_{i})+|g^{\text{target}}(x_{i})\setminus g(x_{i})|-n_{g^{\text{target}}}(x_{i})}{n_{g^{\text{target}}}(x_{i})},1\right)+\frac{1+c}{c\cdot m}\sum_{i\notin A_{g}}\frac{|g^{\text{target}}(x_{i})\setminus g(x_{i})|}{n_{g^{\text{target}}}(x_{i})}$
$\leq$ $\frac{1+c}{m}\sum_{i\in A_{g}}\min\left(\frac{n_{g}(x_{i})}{n_{g^{\text{target}}}(x_{i})}-1,1\right)+\frac{1+c}{m}\sum_{i\in A_{g}}\frac{|g^{\text{target}}(x_{i})\setminus g(x_{i})|}{n_{g^{\text{target}}}(x_{i})}+\frac{1+c}{c\cdot m}\sum_{i\notin A_{g}}\frac{|g^{\text{target}}(x_{i})\setminus g(x_{i})|}{n_{g^{\text{target}}}(x_{i})}$
@@
For any $g$ satisfying $\widehat{\ell}^{\textrm{recall}}(g)\leq\overline{r}$ and $\frac{1}{m}\sum_{i\in[m]}\log\frac{n_{g}(x_{i})\wedge 4n_{g^{\dagger}}(x_{i})}{n_{g^{\dagger}}(x_{i})\wedge 4n_{g}(x_{i})}\leq 6\overline{r}+4\overline{p}+\frac{2}{m}$, we have $\frac{1}{m}\sum_{i\in[m]:\frac{n_{g}(x_{i})}{n_{g^{\text{target}}}(x_{i})}\geq 1}\log\frac{n_{g}(x_{i})\wedge 2n_{g^{\text{target}}}(x_{i})}{n_{g^{\text{target}}}(x_{i})}\leq 18\overline{r}+12\overline{p}+\frac{4}{m}$.
Proof of Lemma 3 Let $B=\{i|\frac{n_{g}(x_{i})}{n_{g^{\text{target}}}(x_{i})}\geq\frac{1}{2}\}$ and $C=\{i|\frac{1}{2}\leq\frac{n_{g^{\dagger}}(x_{i})}{n_{g^{\text{target}}}(x_{i})}\leq 2\}$. For any $i\in B$, the value of $\log\frac{n_{g}(x_{i})\wedge 2n_{g^{\text{target}}}(x_{i})}{n_{g^{\text{target}}}(x_{i})}$ is in [1,1][-1,1]. Then we have
@@
$=$ $\sum_{i\in C\cap B}\log\frac{n_{g}(x_{i})\wedge 4n_{g^{\dagger}}(x_{i})}{n_{g^{\dagger}}(x_{i})\wedge 4n_{g}(x_{i})}+\sum_{i\in C\cap B}\log\frac{n_{g^{\dagger}}(x_{i})\wedge 4n_{g}(x_{i})}{n_{g^{\text{target}}}(x_{i})}+|\neg C|$
$-\sum_{i\in B}\log\frac{n_{g}(x_{i})\wedge n_{g^{\text{target}}}(x_{i})}{n_{g^{\text{target}}}(x_{i})}$
$\leq$ $\sum_{i=1}^{m}\log\frac{n_{g}(x_{i})\wedge 4n_{g^{\dagger}}(x_{i})}{n_{g^{\dagger}}(x_{i})\wedge 4n_{g}(x_{i})}+2|\neg C|+2|\neg B|+\sum_{i\in C\cap B}\log\frac{n_{g^{\dagger}}(x_{i})}{n_{g^{\text{target}}}(x_{i})}+|\neg C|$
$-\sum_{i\in B}\log\frac{n_{g}(x_{i})\wedge n_{g^{\text{target}}}(x_{i})}{n_{g^{\text{target}}}(x_{i})}$
@@
For any hypothesis $g$ with $\widehat{\ell}^{\textrm{recall}}(g)\leq\overline{r}$ and any subset $S\subset[m]$, we have
$\sum_{i\in S\cap B}\log\frac{n_{g}(x_{i})\wedge n_{g^{\text{target}}}(x_{i})}{n_{g^{\text{target}}}(x_{i})}\geq-2m\overline{r}-1\,.$
@@
$\min\sum_{i\in S\cap B}\log\frac{n_{g}(x_{i})\wedge n_{g^{\text{target}}}(x_{i})}{n_{g^{\text{target}}}(x_{i})}$
@@
$\min\sum_{i\in S\cap B}\log\frac{n_{g}(x_{i})\wedge n_{g^{\text{target}}}(x_{i})}{n_{g^{\text{target}}}(x_{i})}\geq-2m\overline{r}-1\,.$
@@
$\sum_{i\in S}\log\frac{n_{g}(x_{i})\wedge n_{g^{\text{target}}}(x_{i})}{n_{g^{\text{target}}}(x_{i})}\geq\min\sum_{i\in S\cap B}\log\frac{n_{g}(x_{i})\wedge n_{g^{\text{target}}}(x_{i})}{n_{g^{\text{target}}}(x_{i})}-|\neg B|\geq-4m\overline{r}-1\,.$
@@
For any hypothesis $g$ with $\widehat{\ell}^{\textrm{precision}}(g)\leq\overline{p}$ and any subset $S\subset[m]$, we have
@@
For any $g$ withe $\widehat{\ell}^{\textrm{recall}}(g)\leq\overline{r}$, we have
@@
For any $g$ withe $\widehat{\ell}^{\textrm{precision}}(g)\leq\overline{p}$, we have
@@
Proof of Lemma 6 When $\frac{n_{g}(x_{i})}{n_{g^{\text{target}}}(x_{i})}<\frac{1}{2}$, we have